# Role of flow topology in wind-driven wildfire propagation

### v1

@@ Abstract
Wildfires propagate through intricate interactions between wind, fuel, and terrain, resulting in complex behaviors that pose challenges for accurate predictions. This study investigates the interaction between wind velocity topology and wildfire spread dynamics, aiming to enhance our understanding of wildfire spread patterns. We revisited the non-dimensionalizion of the governing combustion model by incorporating three distinct time scales. This approach revealed two new non-dimensional numbers, contrasting with the conventional non-dimensionalization that considers only a single time scale. Through scaling analysis, we analytically identified the critical determinants of transient wildfire behavior and established a state-neutral curve, indicating where initial wildfires extinguish for specific combinations of the identified non-dimensional numbers. Subsequently, a wildfire transport solver was developed using a finite difference method, integrating compact schemes and implicit-explicit Runge-Kutta methods. We explored the influence of stable and unstable manifolds in wind velocity on the transport of wildfire under steady wind conditions defined using a saddle-type fixed point flow, emphasizing the role of the non-dimensional numbers. Additionally, we considered the benchmark unsteady double-gyre flow and examined the effect of unsteady wind topology on wildfire propagation, and quantified the wildfire response to varying wind oscillation frequencies and amplitudes using a transfer function approach. The results were also compared to Lagrangian coherent structures (LCS) used to characterize the correspondence of manifolds with wildfire propagation. The comprehensive approach of utilizing the manifolds computed from wind topology provides valuable insights into wildfire dynamics across diverse wind scenarios, offering a potential tool for improved predictive modeling and management strategies.
@@ 1 Department of Mechanical Engineering, University of Utah, Salt Lake City, UT, USA
3Department of Mechanical Engineering, San José State University, San José, CA, USA
4Wildfire Interdisciplinary Research Center (WIRC), San José State University, San José, CA, USA
@@ 1 Introduction
Mathematical models of wildfires provide crucial insights into the complex interactions between the firefront, vegetation, and wind patterns, which collectively constitute the fundamental mechanisms driving wildfire ignition and progression. In a comprehensive review, Sullivan (2009a) detailed the development of physics-based wildfire models, while Sullivan (2009b) explored empirical and quasi-empirical models, with challenges and limitations in accurately simulating fire behavior discussed in Sullivan (2009c) and Or et al. (2023). In the present work, we focus on continuum physics-based wildfire models and specifically the nonlinear convection-diffusion-reaction (CDR) model proposed in Asensio and Ferragut (2002). The mathematical properties of the nonlinear CDR wildfire model, including the existence of traveling wave solutions and uniqueness of solutions, are established in Asensio and Ferragut (2002). Notably, Babak et al. (2009) investigated the profound impact of wind direction on the firefront’s advection within a one-dimensional analytical framework, demonstrating that wind could either accelerate or decelerate fire progression depending on its flow direction while affirming the uniqueness of solutions under varying wind conditions. Similarly, Reisch et al. (2024) investigated the influence of diffusion and reaction parameters, along with wind speed, on wildfire propagation speed, unburned biomass fuel, and maximum achievable fire temperature by examining various simplified sub-models within the wildfire CDR framework.
@@
Numerous studies have explored the influence of wind direction on wildfire behavior.  Fendell and Wolff (2001) introduced an empirical mathematical expression that relates unidirectional wind speed to the fire spread rate for a given fuel load, based on quasi-steady firefront propagation experiments. In related work,  Rossa and Fernandes (2018) developed an empirical model that accounts for wind effects, examining various fuels and their spatial orientations relative to the wind.  Atchley et al. (2021) further advanced these studies by investigating the spatially heterogeneous distribution of fuel, while  Simpson et al. (2013); Li et al. (2021) analyzed the influence of terrain slope on wildfire dynamics. Notably,  Sun et al. (2009) investigated fire-boundary layer interactions through a fully coupled wildfire-atmosphere large eddy simulation for grassland fires. Their study highlighted how a roll-dominated atmospheric boundary layer and its intricate local flow features could influence fire spread and how the fire, in turn, affects the boundary layer.  Filippi et al. (2018) further corroborated the interaction between the fire and boundary layer by performing multi-scale high-resolution large-eddy wildfire simulations, revealing the significance of intricate turbulence effects. Similarly,  Bürger et al. (2020) demonstrated that even minor quasi-temporal variations in wind direction can significantly alter both the propagation direction and the spread length of wildfires.
@@
Despite the growing interest in wind-driven wildfire transport, our understanding of how coherent flow structures in wind influence wildfire propagation patterns remains limited. This paper aims to address this gap by leveraging concepts from dynamical systems theory to identify and predict “coherent patterns” in wind-driven wildfire propagation. In particular, stable and unstable manifolds, identified using Lagrangian coherent structures (LCS), have been utilized in the past to identify templates for convective transport (Shadden et al., 2005; Allshouse and Peacock, 2015; Haller, 2015). LCS has been applied across various fields, including biomedical (Arzani and Shadden, 2012), turbulent (Wilson et al., 2013), environmental (BozorgMagham et al., 2013), and reactive flows (Mahoney et al., 2012). Within the context of wildfires, recent studies from the atmospheric science community have employed the LCS framework to qualitatively predict the evolution of smoke plumes from wildfire events (Curbelo and Rypina, 2023; Allen et al., 2024; Jarvis et al., 2024). However, the direct relevance of LCS to wildfire propagation has yet to be explored. Given that wildfire transport is not purely advective, it remains unclear how flow topology can be utilized and under what conditions its role becomes significant. Gaining such fundamental insights is crucial, as it enables the prediction of wildfire growth patterns without the need to solve computationally expensive transport problems, assuming the wind velocity field is known. The attracting LCS (unstable manifolds) tend to draw transported material, such as fire lines, towards themselves, which could indicate regions of heightened risk. Conversely, the repelling LCS (stable manifolds) may potentially push fire lines away, thereby creating safer zones for positioning firefighters.
@@
In this paper, we consider the Asenio wildfire model (Asensio and Ferragut, 2002), which incorporates key pathways in wildfire transport, including wind-driven convection, nonlinear diffusion due to heat and radiation, fuel-limited reaction, and simplified natural convection. The model, however, does not account for several critical factors, such as terrain slope (Coen, 2018) and firebrands (Tohidi and Kaye, 2017). Additionally, we assume the wind velocity is known and provided, thereby neglecting the heating-induced buoyancy forces that can alter wind velocity—factors that have been incorporated in more complex coupled weather–wildland models (Mandel et al., 2011; Coen et al., 2013; Coen, 2015). While we acknowledge the relative simplicity of the Asenio model, especially given the multiscale nature of wildfires (Hädrich et al., 2021; Hudson et al., 2020), this simplicity allows us to focus on the fundamental factors and understand the role of flow topology in the interplay between key physical processes (reported using nondimensional groups). From a fundamental perspective, we approach the problem by modeling nonlinear convection-diffusion-reaction transport, where limited fuel (vegetation) acts as a source (reaction term) sustaining the high-temperature fire spread. This approach prioritizes a fundamental reactive flow mindset over a realistic, wildfire-specific modeling approach. The precise flame and combustion chemistry are not modeled, as these phenomena occur on scales much smaller than those at which wind velocity data is typically measured or modeled. Nonetheless, understanding the behavior of this particular class of reacting flows will lay the groundwork for developing more complex models in the future.
@@
The paper is organized as follows: In Section 2, we revisit the wildfire spread model presented by Asensio and Ferragut (2002) and proceed to non-dimensionalize it, introducing previously unrecognized dimensionless numbers. This section also details the development of a wildfire transport solver based on finite difference methods. Section 3 presents a scaling analysis to identify the key factors influencing wildfire behavior. In Section 4, we explore the characteristics of transient wildfire dynamics under steady wind conditions with saddle-type fixed points and analyze the impact of unsteady wind velocities modeled by a double-gyre flow.
@@ 2.1 Mathematical wildfire model
The mathematical nonlinear convection-diffusion-reaction model for wildland fire propagation, proposed by Asensio and Ferragut (2002), is based on the principles of energy and species conservation, considering the quantity of fuel while disregarding its specific intricate composition. The reaction rate $(r)$ in this model is characterized by the Arrhenius equation, $r=A\exp{\left(-E_{\mathrm{A}}/(RT)\right)}$, and the fuel burning rate $(S_{Y})$ is proportional to this reaction. In wildfire scenarios, sub-grid radiation flux through fuel parcels is an important mode of heat transfer, characterized by the Stefan-Boltzmann law, $q_{\text{rad}}=4\sigma\delta T^{3}$. Furthermore, natural convection is encapsulated by Newton’s law of cooling, $q_{\text{conv}}=h(T-T_{\infty})$. Integrating these components–convection, radiation, reaction, and natural convection–the proposed one-way coupled model for flame propagation and fuel burning is mathematically formulated as
@@
where $\rho$ is the density of air, $C$ is the constant pressure specific heat, $T$ is the absolute temperature, $\overrightarrow{\mathbf{v}}$ is the velocity vector, $\sigma$ stands for the Stefan-Boltzmann constant, $\delta$ represents the optical path length for radiation, $k$ indicates thermal conductivity, $s(T)^{+}$ characterizes the dimensional phase change function, $Y$ is the fuel mass fraction, $H$ represents the heat of combustion, $r$ is the reaction rate, $h$ is the natural heat convection coefficient, $T_{\infty}$ stands for the ambient temperature, $A$ represents the pre-exponential factor for a first order kinetic scheme, $E_{\mathrm{A}}$ is the activation energy, and $R$ is the universal gas constant.
The mathematical model was originally non-dimensionalized using the Frank-Kamenetskii change of variables method, considering only one characteristic temporal scale ($t_{0}$) and spatial scale ($l_{0}$). It involves applying the following change of variables to non-dimensionalize the wildfire equation.
@@
where $\xi$ and $\eta$ represent the normalized spatial lengths along the $x$ and $y$ directions, respectively, using the characteristic length scale $l_{0}=\sqrt{(t_{0}k)/(\rho C)}$. $\tau$ denotes the dimensionless evolution time normalized using the characteristic temporal scale $t_{0}=(\mathrm{e}^{(1/\epsilon)}\epsilon)/(qA)$, where $\epsilon=(RT_{\infty})/(E_{\mathrm{A}})$ represents the inverse activation energy and $q=(HY_{0})/(CT_{\infty})$ is the non-dimensional reaction heat. In addition, $\overrightarrow{\mathbf{w}}$ denotes the normalized velocity, and $\beta$ represents the normalized fuel mass fraction. Using the non-dimensional absolute temperature ($\overline{T}$) and non-dimensional phase temperature ($\overline{T}_{\mathrm{pc}}$), the non-dimensional phase function is defined as
@@ 2.2 New non-dimensional wildfire model
In the wildfire model described by Eq. 1, three distinct characteristic temporal scales are evident: fuel reaction ($t_{r}$), flow convection ($t_{f}$), and diffusion ($t_{d}$). However, Asensio and Ferragut (2002) assumed these temporal scales to be identical when using the Frank-Kamenetskii method to derive the non-dimensionalized form of the wildfire equation shown in Eq. 4. Here, we define the three distinct characteristic temporal scales for the different processes according to Frank-Kamenetskii theory, along with the wind velocity:
@@
[…]

### v2

@@ Abstract
Wildfires propagate through interactions between wind, fuel, and terrain, resulting in complex behaviors that challenge accurate predictions. This study investigates the interaction between wind velocity topology and wildfire dynamics, aiming to enhance our understanding of wildfire spread patterns through a simplified nonlinear convection-diffusion-reaction wildfire model, adopting a fundamental reactive flow dynamics perspective. We revisited the non-dimensionalizion of the governing combustion model by incorporating three distinct time scales. This approach revealed two new non-dimensional numbers, contrasting with the conventional non-dimensionalization that considers only a single time scale. Through scaling analysis, we analytically identified the critical determinants of transient wildfire behavior and established a state-neutral curve, indicating where initial wildfires extinguish for specific combinations of the identified non-dimensional numbers. Subsequently, a wildfire transport solver was developed, integrating upwind compact schemes and implicit-explicit Runge-Kutta methods. We explored the influence of stable and unstable manifolds in wind topology on the transport of wildfire under steady wind conditions defined using a saddle-type fixed point flow, emphasizing the role of the non-dimensional numbers. Additionally, we considered the benchmark unsteady double-gyre flow and examined the effect of unsteady wind topology on wildfire propagation, and quantified the wildfire response to varying wind oscillation frequencies and amplitudes using a transfer function approach. The results were compared to Lagrangian coherent structures (LCS) used to characterize the correspondence of manifolds with wildfire propagation. The approach of utilizing the wind flow manifolds provides valuable insights into wildfire dynamics across diverse wind scenarios, offering a potential tool for improved predictive modeling and management strategies.
@@ 1 Department of Mechanical Engineering, University of Utah, Salt Lake City, UT, USA
3Department of Fire Protection Engineering, University of Maryland, College Park, MD, USA
4NSF-IUCRC Wildfire Interdisciplinary Research Center (WIRC), San José, CA, USA
@@ 1 Introduction
Mathematical models of wildfires provide crucial insights into the complex interactions between the firefront, vegetation, and wind patterns, which collectively constitute the fundamental mechanisms driving wildfire ignition and progression. In a comprehensive review, Sullivan (2009a) detailed the development of physics-based wildfire models, while Sullivan (2009b) explored empirical and quasi-empirical models, with challenges and limitations in accurately simulating fire behavior discussed in Sullivan (2009c) and Or et al. (2023). While the simplest empirical models establish functional relationships between the fire spread rate and key parameters—such as slope, wind speed, fuel type, and bulk density—the other end of the spectrum involves the use of three-dimensional coupled computational fluid dynamics-wildfire (CFD) solvers, which aim to comprehensively capture physical and chemical interactions across a broad range of scales. Positioned between these empirical functional relations and CFD tools, several continuum physics-based wildfire models—proposed over the years by Asensio and Ferragut (2002); Margerit and Séro-Guillaume (2002); Coen et al. (2007); Mandel et al. (2008); Simeoni et al. (2011); Bürger et al. (2020); Vogiatzoglou et al. (2024); Navas-Montilla et al. (2024) —aim to describe the fundamental reactive flow physics, specifically through the nonlinear convection-diffusion-reaction (CDR) framework, while maintaining computational tractability. In this approach, the modeling involves a simplified description of combustion kinetics and thermal energy transfer, where all the physical quantities are averaged over the local plantation height, resulting in a computationally efficient system of coupled differential equations that describes the spatiotemporal evolution of the two-dimensional fields of temperature and combustibles. Furthermore,  Gollner et al. (2015); Verma et al. (2022) remarked that these physics-based CDR wildfire models effectively resolve the “flame” scales, as the pyrolysis processes occurring at the “vegetation” scale are not captured by these reaction models.
@@
Numerous studies have explored the influence of wind direction on wildfire behavior within the contexts of empirical-based, physics-based CDR, and fully coupled CFD methodologies. In the realm of empirical-based wildfire models, Babak et al. (2009) investigated the profound impact of wind direction on the firefront’s advection within a one-dimensional analytical framework, demonstrating that wind could either accelerate or decelerate fire progression depending on its flow direction while affirming the uniqueness of solutions under varying wind conditions. Similarly,  Fendell and Wolff (2001) introduced an empirical mathematical expression that relates unidirectional wind speed to the fire spread rate for a given fuel load, based on quasi-steady firefront propagation experiments. In related work, Rossa and Fernandes (2018) developed an empirical model that accounts for wind effects, examining various fuels and their spatial orientations relative to the wind. Atchley et al. (2021) further advanced these studies by investigating the spatially heterogeneous distribution of fuel, while  Simpson et al. (2013); Li et al. (2021); Bi et al. (2022, 2023), along with the references cited therein, analyzed the influence of terrain slope on wildfire dynamics.
Within the CDR wildfire framework, Asensio and Ferragut (2002) studied the significance of flow advection on firefronts driven by moderate, unidirectional wind over both homogeneous and heterogeneous fuel beds. Coen et al. (2007) demonstrated that these physics-based wildfire models efficiently capture the temporal temperature data sampled from real-time weather measurements, thereby suggesting their applicability in short-range wildfire behavior forecasting. In their subsequent work, Mandel et al. (2008) employed data assimilation methodologies to calibrate parameters within the CDR model and improve its prediction accuracy. Simeoni et al. (2011) studied the reliability of these models by qualitatively reproducing fire behavior observed in field experiments.  Bürger et al. (2020) highlighted that even minor quasi-temporal variations in wind direction and firebreaks can significantly alter both the propagation direction and the spread rate of wildfires. Furthermore, Navas-Montilla et al. (2024) provided theoretical and numerical insights on wildfire advection under the influence of moisture content, heterogeneous fuel distribution, and non-flat topography. Reisch et al. (2024) investigated the influence of diffusion and reaction parameters, along with wind speed, on wildfire propagation speed, unburned biomass fuel, and maximum achievable fire temperature by examining various simplified sub-models.
Utilizing fully coupled CFD methodologies,  Sun et al. (2009) investigated fire-boundary layer interactions through a fully coupled atmosphere-wildfire large eddy simulation for grassland fires. Their study highlighted how a roll-dominated atmospheric boundary layer and its intricate local flow features could influence fire spread and how the fire, in turn, affects the boundary layer. Filippi et al. (2018) further corroborated the interaction between the fire and boundary layer by performing multi-scale high-fidelity large-eddy wildfire simulations, revealing the significance of intricate turbulence effects. The studies by Clark et al. (1996); Coen and Riggan (2014); Coen and Schroeder (2015) utilize instantaneous transient wind velocity vectors to investigate their correspondence with the advection dynamics of wildfires over realistic terrain and fuel distributions, employing the fully coupled atmosphere (weather)–wildland fire model simulations. Their findings indicate that fire progression is predominantly governed by the wind velocity vector patterns in the vicinity of the fire location.
@@
Given the significant growing interest in understanding wind-driven wildfire transport through the wind velocity fields, the present work takes a different approach to understand how coherent flow structures in wind influence wildfire propagation patterns, rather than relying on local wind velocity vectors. This paper aims to leverage concepts from dynamical systems theory to identify and predict “coherent patterns” in wind-driven wildfire propagation. In particular, stable and unstable manifolds, identified using Lagrangian coherent structures (LCS), have been utilized in the past to identify templates for convective transport (Shadden et al., 2005; Allshouse and Peacock, 2015; Haller, 2015). LCS has been applied across various fields, including biomedical (Arzani and Shadden, 2012), turbulent (Wilson et al., 2013), environmental (BozorgMagham et al., 2013), and reactive flows (Mahoney et al., 2012). Within the context of wildfires, recent studies from the atmospheric science community have employed the LCS framework to qualitatively predict the evolution of smoke plumes from wildfire events (Curbelo and Rypina, 2023; Allen et al., 2024; Jarvis et al., 2024). However, the direct relevance of LCS to wildfire propagation has yet to be explored. Given that wildfire transport is not purely advective, it remains unclear how flow topology can be utilized and under what conditions its role becomes significant. Gaining such fundamental insights is crucial, as it enables the prediction of wildfire spread patterns without the need to solve computationally expensive transport problems, assuming the wind velocity field is known. The attracting LCS (unstable manifolds) tend to draw transported material, such as fire lines, towards themselves, which could indicate regions of heightened risk. Conversely, the repelling LCS (stable manifolds) may potentially push fire lines away, thereby creating safer zones for positioning firefighters.
@@
In this paper, we consider the Asensio CDR wildfire model (Asensio and Ferragut, 2002), which incorporates key pathways in wildfire transport, including wind-driven convection, nonlinear diffusion due to heat and radiation, fuel-limited reaction, and simplified natural convection. The model, however, does not account for several critical factors, such as firebrands (Tohidi and Kaye, 2017). Additionally, we assume the wind velocity is known and provided, thereby neglecting the heating-induced buoyancy forces that can alter wind velocity—factors that have been incorporated in more complex coupled weather–wildland models (Mandel et al., 2011; Coen et al., 2013; Coen, 2015). While we acknowledge the relative simplicity of the Asensio model, especially given the multiscale nature of wildfires (Hädrich et al., 2021; Hudson et al., 2020), this simplicity allows us to focus on the fundamental factors and understand the role of flow topology in the interplay between key physical processes (reported using nondimensional groups). From a fundamental perspective, we approach the problem by modeling nonlinear convection-diffusion-reaction transport, where limited fuel (vegetation) acts as a source (reaction term) sustaining the high-temperature fire spread. This approach prioritizes a fundamental reactive flow mindset over a realistic, wildfire-specific modeling approach. The precise flame and combustion chemistry are not modeled within these typical CDR wildfire models, as these phenomena occur on vegetation scales much smaller than those at which wind velocity data is typically measured or modeled. Nonetheless, understanding the behavior of this particular class of reactive flows will lay the groundwork for developing more complex wildfire models in the future.
@@
The paper is organized as follows: In Section 2, we revisit the wildfire model presented by Asensio and Ferragut (2002) and proceed to non-dimensionalize it, introducing previously unrecognized dimensionless numbers. This section also details the development of a wildfire transport solver based on finite difference methods. Section 3 presents a scaling analysis to identify the key factors influencing wildfire behavior within the physics-based CDR framework. In Section 4, we explore the characteristics of transient wildfire dynamics under steady wind conditions with saddle-type fixed points and analyze the impact of unsteady wind velocities modeled by a double-gyre flow.
@@ 2.1 Mathematical wildfire model
The mathematical nonlinear convection-diffusion-reaction model for wildland fire propagation, proposed by Asensio and Ferragut (2002), is based on the principles of energy and species conservation, considering the quantity of fuel while disregarding its specific intricate composition. The reaction rate $(r)$ in this model is characterized by the Arrhenius equation, $r=A\exp{\left(-E_{\mathrm{A}}/(RT)\right)}$, and the fuel burning rate $(S_{Y})$ is proportional to this reaction. In wildfire scenarios, sub-grid radiation flux through fuel parcels is an important mode of heat transfer, characterized by the Stefan-Boltzmann law, $q_{\text{rad}}=4\sigma\delta T^{3}$, adopting the local radiative model (Rosseland, 1931). Furthermore, natural convection is encapsulated by Newton’s law of cooling, $q_{\text{conv}}=h(T-T_{\infty})$. Integrating these components–convection, radiation, reaction, and natural convection–the proposed one-way coupled model for flame propagation and fuel burning is mathematically formulated as
@@
where $\rho$ is the density, $C$ is the constant pressure specific heat, $T$ is the absolute temperature, $\overrightarrow{\mathbf{v}}$ is the velocity vector, $\sigma$ stands for the Stefan-Boltzmann constant, $\delta$ represents the optical path length for radiation, $k$ indicates thermal conductivity, $s(T)^{+}$ characterizes the dimensional phase change function, $Y$ is the fuel mass fraction —the ratio of the mass to the initial mass, $H$ represents the heat of combustion, $r$ is the reaction rate, $h$ is the natural heat convection coefficient, $T_{\infty}$ stands for the ambient temperature, $A$ represents the pre-exponential factor for a first order kinetic scheme, $E_{\mathrm{A}}$ is the activation energy, and $R$ is the universal gas constant. It should be noted that all physical variables in Eq. 1 are spatially averaged quantities, with the averaging performed over the fuel height, perpendicular to the ground, accounting for both air and fuel properties. Furthermore, the influence of topography can be incorporated into the wildfire model by redefining the velocity vector $\overrightarrow{\mathbf{v}}$ as proportional to the topographic gradient, as shown in Nelson Jr (2002), with an additional advection correction factor as described in Grasso and Innocente (2020); Nieding et al. (2024); Navas-Montilla et al. (2024).
The mathematical model was originally non-dimensionalized using the Frank-Kamenetskii change of variables method, considering only one characteristic temporal scale ($t_{0}$) and spatial flame scale ($l_{0}$). It involves applying the following change of variables to non-dimensionalize the wildfire equation.
@@
where $\xi$ and $\eta$ represent the normalized spatial lengths along the $x$ and $y$ directions, respectively, using the characteristic length scale $l_{0}=\sqrt{(t_{0}k)/(\rho C)}$. $\tau$ denotes the dimensionless evolution time normalized using the characteristic temporal scale $t_{0}=(\mathrm{e}^{(1/\epsilon)}\epsilon)/(qA)$, where $\epsilon=(RT_{\infty})/E_{\mathrm{A}}$ represents the inverse activation energy and $q=(HY_{0})/(CT_{\infty})$ is the non-dimensional reaction heat. In addition, $\overrightarrow{\mathbf{w}}$ denotes the normalized velocity, and $\beta$ represents the normalized fuel mass fraction. Using the non-dimensional absolute temperature ($\overline{T}$) and non-dimensional phase temperature ($\overline{T}_{\mathrm{pc}}$), the non-dimensional phase function is defined as
@@ 2.2 New non-dimensional wildfire model
In the wildfire model described by Eq. 1, three distinct characteristic temporal scales are evident: fuel reaction ($t_{r}$), flow convection ($t_{f}$), and diffusion ($t_{d}$). However, Asensio and Ferragut (2002) assumed these temporal scales to be identical when using the Frank-Kamenetskii method to derive the non-dimensionalized form of the wildfire equation shown in Eq. 4. Here, we define the three distinct characteristic temporal scales for the three different processes according to Frank-Kamenetskii theory, along with the wind velocity:
@@
In this context, the Damköhler number $\textit{Da}=t_{d}/t_{r}=(\rho Cl_{0}^{2}qA)/(\epsilon e^{(1/\epsilon)}k)$ represents the ratio of the diffusion time scale to the reaction time scale, indicating how fast the reaction rate is in comparison to the diffusion rate. Similarly, Peclet number $\textit{Pe}=t_{d}/t_{f}=(\rho CU_{\infty}l_{0})/k$ signifies the ratio of the diffusion time scale to the flow convection time scale, elucidating the relationship between convection and diffusion. Furthermore, a new non-dimensional number, $\Phi=t_{f}/t_{r}=\textit{Da}/\textit{Pe}$, has been identified and defined as the ratio of flow convection rate to reaction rate, thus establishing a relationship between the Damköhler number and the Peclet number. This process of non-dimensionalization has unveiled a previously unnoticed new non-dimensional number ($\Phi$) and Damköhler number inherent within the wildfire combustion model under consideration. It is imperative to underscore that the resulting new non-dimensional combustion model (Eq. 8) reduces to the conventional form represented by Eq. 4 in instances where both Da and $\Phi$ values converge to 1.
Based on reported data in the literature (Fendell and Wolff, 2001; Asensio and Ferragut, 2002; Cheney et al., 2012; Amini et al., 2019; Bürger et al., 2020; Verma et al., 2022), the following variables govern the two identified non-dimensional numbers: the ambient wind speed $U_{\infty}$, ranging from $10^{-3}\,\text{m}\,\text{s}^{-1}$ (no wind) to $25\,\text{m}\,\text{s}^{-1}$; the flame length scale $l_{0}$, which varies from $\mathcal{O}(10^{-1})$ to $\mathcal{O}(10^{2})\,\text{m}$; and the pre-exponential factor $A$, spanning $10^{6}$ to $10^{14}\,\text{s}^{-1}$, with higher values observed for dead fuels. These ranges were used to guide the selection of the non-dimensional numbers within our non-dimensional equation. It is important to note a limitation within the physics-based CDR wildfire model, where the variation of non-dimensional numbers are “shaped” by other spatially averaged physical variables accounting for both fuel and air properties ($\rho=100\,\text{kg}\,\text{m}^{-3}$, $C=1\,\text{kJ}\,\text{kg}^{-1}\,\text{K}^{-1}$, $k=1\,\text{W}\,\text{m}^{-1}\,\text{K}^{-1}$, $T_{\infty}=300\,\text{K}$). As pointed out by Vogiatzoglou et al. (2024), the spatial averaging process over plantation height and the refinement of the choice of mean values would constitute an independent fluid mechanics problem, which has been left for future studies aimed at improving wildfire CDR models.
For the same choice of mean values, the Da number is found to vary from $\mathcal{O}(1)$ to $\mathcal{O}(10^{7})$ considering extreme values of $l_{0}$ and $A$ values. On the other hand, the $\Phi$ value, representing the convection term influence, ranges from $\mathcal{O}(10^{2})$ to $\mathcal{O}(10^{-7})$, computed based on the chosen range of $l_{0}$, $A$, and $U_{\infty}$ values. In this study, the newly derived non-dimensional wildfire equation is employed to investigate fire propagation under various wind conditions, with Da ranging from $1$ to $10^{5}$, characterizing the strength of diffusion, and the wind topologies associated with $\Phi$ values ranging from $10^{2}$ to $10^{-3}$, covering the wind regime from very calm to moderate. Furthermore, the effect of the reaction process is investigated by varying the parameter $\epsilon$ between $0.03$ and $0.045$, corresponding to different activation energy values. It is anticipated that in scenarios with significantly stronger winds, the influence of convection and wind topology on wildfire propagation will become more pronounced.